the length of a rectangular board is 9 inches more than its width, the area is 112 square inches. Find the measures of the length and the width

Answer by ikleyn(52778)

.
The simplest solution is this:

    Factor 112 into the product of two numbers that differ in 9:


    112 = 56*2 = 28*4 = 14*8 = 7*16.


    7 and 16 are the two appropriate numbers.

Answer. The length is 16 inches, the width is 7 inches.


I am sure other tutors will show you other possible solutions.